OpenStudy (anonymous):

Find the equation of the line that fits the description. Passes through (6,−1) and has zero slope.

OpenStudy (anonymous):

In what form? Point slope? Slope intercept?

OpenStudy (alexandervonhumboldt2):

ok so if slope is 0 the line will have form y=a. lines with 0 slope is horizontal. 

OpenStudy (anonymous):

@AlexandervonHumboldt2 it says equation.

OpenStudy (alexandervonhumboldt2):

i know

OpenStudy (alexandervonhumboldt2):

Passes through (6,−1) and has zero slope. y=mx+b. m=0 substitute your point: -1=6*0+b b=-1

OpenStudy (alexandervonhumboldt2):

b is y-intercept, thus the equation will look like y=-1

pooja195 (pooja195):

Please do not give the answer.

OpenStudy (anonymous):

so when it asks for the equation of a line would i give an answer of -1=6(0)+b or I am confused

pooja195 (pooja195):

We can do this another way i am confused by the way alex did this... I personally perfer the point slope method either way we get the same answer \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Can you plug in the info the formula?

OpenStudy (alexandervonhumboldt2):

what confused way? we have that slope =0 and y intercept is -1. y=mx+b y=0*x-1 y=-1

OpenStudy (anonymous):

y-1=0(x-6)

OpenStudy (anonymous):

then add 1 both sides and y=1

pooja195 (pooja195):

No first distribute the 0

OpenStudy (anonymous):

distributing 0 is 0 right?

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

so its y-1=0 then you add 1 to both sides and y=1

pooja195 (pooja195):

You didnt plug in a coordinate right

pooja195 (pooja195):

y+1=0(x-6) Remeber that a neg + neg= pos

pooja195 (pooja195):

And now what will your answer be y=?

OpenStudy (anonymous):

y=-1

pooja195 (pooja195):

Correct well done :)

OpenStudy (anonymous):

so when it asks for a equation how do i put it?

pooja195 (pooja195):

y=-1 is your equation its a horizontal line

OpenStudy (anonymous):

oh okay thanks!!

OpenStudy (anonymous):

:)

pooja195 (pooja195):

No problem =)

OpenStudy (athecool):

Hey pooja195 would you help me at http://openstudy.com/study#/updates/565cf04ee4b0603f58e8dce6?

OpenStudy (athecool):

PLease???